OpenStudy (anonymous):

Ten minus five b equals negative eight minus fourteen b... A. 2/19 B. 18/19 C. -2/19 D. -2

OpenStudy (anonymous):

10-5b=8-14b

OpenStudy (anonymous):

10-5b = -8-14b -5b +14b = -8 - 10 b = ?

OpenStudy (anonymous):

combine the like terms together